. Following information is available for Companies Ace Ltd. and Pace Ltd.: (₹ in lacs) Particulars Ace Ltd. Pace Ltd. Long term Debt 625 700 Equity 2100 2850 Current assets 450 550 Current liabilities 300 375 Net Profit 115 178 Revenue (net) 355 452 a. Compute Debt-equity ratio, current ratio for both companies


1
Expert's answer
2022-03-04T13:19:02-0500

Ace ltd

Debt Equity Ratio= "\frac{Total liabilities}{Share Equity}"

"\frac{625+300}{2100}= 0.44"


"Current Ratio= \frac{Current Assets}{Current Liabilities}"

"= \frac{450}{300}= 1.5"

Pace ltd

"Debt Equity Ratio= \frac{Total liabilities}{Share Equity}"

"\frac{700+375}{2850}= 0.44"

"Current Ratio= \frac{Current Assets}{Current Liabilities}"

"= \frac{550}{375}= 1.47"